Abstract

Let t be an integer and suppose G is a graph on n vertices. Given an ordered set of t vertices in G and a set of t distances, we find a sharp degree condition for G to contain a hamiltonian cycle with the t vertices in order with approximately the corresponding prescribed distances between them. We also show similar results for partitioning the graph into paths between chosen vertices or paritioning into cycles containing chosen vertices.

Advisor: Ron Gould